Solve,show your work to help me understand.
An algebra book weighs 6oz less than twice as much as a grammer book. if 5 algebra books weigh the same as 8 grammer books, how much does an algebra book weigh?

Answer by robertb(4012)

a = 2g-6,
5a = 8g.
Then, by substitution, 5(2g-6) = 8g,
10g-30 = 8g,
10g-8g = 30,
2g = 30,
g = 15. A grammar book weighs 15 ounces, an algebra book, 2*15 - 6 = 24 ounces.